OpenStudy (anonymous):

Karen uses her credit card to purchase a new television for $695.20. She can pay off up to $325 per month. The card has an annual rate of 17.9% compounded monthly. How much will she pay in interest? $6.28 $16.96 $20.44 $62.16

OpenStudy (tkhunny):

Just do it one month at a time. Shall we assume payments are at the END of the month? :-)

OpenStudy (anonymous):

yes :)

OpenStudy (tkhunny):

Let's see it. Month 1: 695.20*(1 + 0.179/12) - 325.00 = Remaining Balance Month 2: You tell me.

OpenStudy (anonymous):

380.57?

OpenStudy (amistre64):

since the period of payments are each month, and the interest rate is stated in a period of 1 year; let r = interest rate/12 m1 = B(1+r) - P m2 = m1(1+r) - P m3 = m2(1+r) - P ... or written another way m1 = B(1+r) - P m2 = (B(1+r) - P)(1+r) - P m3 = ((B(1+r) - P)(1+r) - P)(1+r) - P ..... or to clean it up even more m1 = B(1+r)^1 - P(1+r)^0 m2 = B(1+r)^2 - P( (1+r)^0 + (1+r)^1 ) m3 = B(1+r)^3 - P( (1+r)^0 + (1+r)^1 + (1+r)^2) let 1+r = k for a clean up, the reccursion becomes \[m_n=B(k)^n-P(1+k+k^2+k^3+k^4+...+k^{n-1})\]

OpenStudy (amistre64):

that 1+k+k^2+k^3 .... is a geometric sum that can be state as (1-k^n)/(1-k) \[m_n=B(k)^n-P\frac{1-k^n}{1-k}\] if we want the nth month such that the final balance is zero \[0=B(k)^n-P\frac{1-k^n}{1-k}\] \[B(k)^n=P\frac{1-k^n}{1-k}\] \[\frac{B(1-k)}{P}(k)^n=1-k^n\] \[\frac{B(1-k)}{P}=k^{-n}-1\] \[\frac{B(1-k)+P}{P}=k^{-n}\] \[\log_k(\frac{P}{B(1-k)+P})=n\] that looks like fun

OpenStudy (amistre64):

the good news is that it looks like its only gonna take about 2.19 months to pay off

OpenStudy (amistre64):

she ends up paying 718.12 for the thing -695.20 -------- 22.92 which prolly has some error in it, but is close enough :)

OpenStudy (anonymous):

that's really confusing 0_o

OpenStudy (amistre64):

lol, when i cant remember a formula, i tend to have to reinvent it

OpenStudy (amistre64):

what formulas do they give you in your material to be able to work with this?

OpenStudy (tkhunny):

Month 1: 695.20*(1 + 0.179/12) - 325.00 = 380.57 Month 2: 380.57*(1 + 0.179/12) - 325.00 = 61.25 Month 3: 61.25*(1 + 0.179/12) - 62.16 = 0.00 It does not take 2.9 months. It takes 3 months. Total Paid: 325.00 + 325.00 + 62.16 = 712.16 Total Interest Paid: 712.16 - 695.20 = 16.96

OpenStudy (amistre64):

yeah, i got to 2.19 and forgot to round the months to 3 :)